​a heap of rice is in the form of cone of diameter 9m and height 3.5m. find the volume of the rice. how much canvas cloth is required to just cover the heap?

Grade:10

1 Answers

Arun
6 years ago
Dear Divya
 
Volume of cone = 1/3 * pie * r * r * h

D = 9 m , r = 9 / 2 m

Volume = (1 / 3) * (22 / 7) * (9/2)*(9/2)*(3.5) = 74.25 m3

Just to cover , the heap, we need the CSAof cone

Slant height of cone = root [ r2 + h2 ] = root [ (9 / 2) * (9 / 2) + (3.5)*(3.5) ] = 5.7008 m

CSA= pie * r * l = (22 / 7) * (9/2) * (5.7008) = 80.6256 m

Cheers!
 
 
Regards
Arun (askIITians forum expert)